—Estimate the net charge of a Glu–Tyr dipeptide at pH 6.0.


1
Expert's answer
2021-09-29T05:32:02-0400

Find the pKa values for Glu

pK1 (alpha carboxylic group) - 2.19, which is less than pH = 6, hence, the alpha carboxylic group will hold a negative (-1) charge.

pK2 (Alpha amino group) - 9.67, which is more than pH = 6, hence it will have a positive charge (+1).

pK3 (side chain) - 4.25, which is less than pH = 6, hence, the side chain carboxylic group will hold a negative (-1) charge.

Now, the pKa values for Tyr are -

pK1 (alpha carboxylic group) - 2.20, which is less than pH = 6, hence, the alpha carboxylic group will hold a negative (-1) charge.

pK2 (Alpha amino group) - 9.11, which is more than pH = 6, hence it will have a positive charge (+1).

Hence, the net charge of the dipeptide Glu-Tyr will be -

(-1) + (+1) + (-1) + (-1) + (+1) = -1.